Simon Compenys year-end balance sheets follow Accounts receivable, net Merchandise inventory Prepeid expenses Plant assets, net $ 32,544 $ 38,041 $ 4e,e38 89,200 62,800 3,588 113,000 85,8 59,800 ,480 9,96 4,448 317.629 290,191 255,622 $562,853 5485,218 5412,600 Liabilities and tquity Accounts payable Long-tern notes payable secured by $137,347 8,362 53,374 sortgages on plant assets Common stock, $18 par value Retained earnings Totel liabilities and equity 159-386132,88187,365 $562,853 $485,218 $412,688 The companys income statements for the years ended December 31, 2017 and 2016, follow Assume that all sales are on $731,789 $577,489 Cost of goods sold Other operating expenses Interest expense $446,342 226,830 146,884 13,280 Total costs and expenses Net incone 36,586 52.25 Earnings per sharye (1) Compute days sales uncollected Days-Days Sales Days Sale (2) Compute accounts recelvable turnover (3) Compute inventory turnover Inventory tunover 2017 4) Compute days sales in inventory DaysDays Sales In
Answer #1
Days Sales uncollected
Year Average Receivables / Sales revenue    * Days = Days sales uncollected
2017 75600 731709 365 37.71 days
2016 57750 577409 365 36.51 days
Note:
Average Accounts receivable = Beginning AR + Ending AR /2
2017 = (89200+62000) /2 = 75600
2016 = (62000+53500) / 2 = 57750
Accounts Receivable turnover
Year Net sales       / Average Accounts receivable = Accounts receivable Turnover
2017 731709 75600 9.68 times
2016 577409 57750 10 times
Inventory Turnover:
Year Cost of goods sold        / Average inventory = Inventory Turnover
2017 446342 99000 4.51 times
2016 375316 72000 5.21 times
Note:
Average inventory = (beginning Inventory+ Ending inventory) /2
2017 = (85000+13000)/2 = 99000
2016 (59000+85000) /2 = 72000
Days sales in inventory
Year Average inventory / Cost of good s sold * Days   = Days sales in Ivnentory
2017 99000 446342 365 80.96 days
2016 72000 375316 365 70.02 days